A merchant blends tea that sells for $5.00 a pound with tea that sells for $2.50 a pound to produce 50 lb of a mixture that sells for $3.00 a pound. How many pounds of each type of tea does the merchant use in the blend?
Answer by ewatrrr(24785)


Hi
produce 50 lb of a mixture
Let x and (50-x)represent the amount of the 5$ and $2.50 mixtures respectively
Question states*** Mixing
5x + 2.50(50-x) =3*50
Solving for x
2.50x = .50*50
x = .50*50/2.50
x = 10lb, the amount of the $5/lb tea. 40lb of $2.50/lb tea.
CHECKING our Answer***
50 + 100 = 150